In a container, there are dimes and nickels. There are 10 more nickels than dimes and the nickels and dimes add up to $3.50. How many nickels and dimes are there?
Answer by Alan3354(69443)

In a container, there are dimes and nickels. There are 10 more nickels than dimes and the nickels and dimes add up to $3.50. How many nickels and dimes are there?
-----------
Here's one way:
n = d + 10
5n + 10d = 350
====================
Here's another:
Without the 10 more nickels, the total is $3.00.
Each set of 1 dime & 1 nickel = 15 cents.
300/15 = 20 sets
--> 20 dimes & 30 nickels